Why the Theory of Turing-Machine Functionalism Works

Marc Horowitz April 1, 1992

If the computational theory of cognition is correct, the most general model we can choose is a Turing machine. Since a Turing machine operates on discrete states and inputs, then there must be an internal representation of mental states, a representation of external stimuli, and the output of the machine must represent some action. The theory of Turingmachine functionalism satis�es these requirements and stands up well to objections.

It is fairly obvious that what is important about thinking is not the details of the input or the output. We should not be concerned with how vision or hearing works, or how muscular contraction works. Similarly, the chemistry of the brain is unimportant. What is important is how stimulus, beliefs, and desires bring about new beliefs, desires, and actions. Turing-machine functionalism puts forth a theory where we abstract out the physical details of stimuli, belief, and desires into intentional states which are independent of the individual, or even the species. According to Block, �what gives internal symbols their meanings is how they function.� How they function is determined by how they affect the output of the Turing machine which represents the cognitive functions of the brain.

This can be compared to behaviorism, which states that �mental states were identical with sets of actual and counterfactual overt behaviors and that inner states of the subject, though no doubt causally implicated in such behaviors, were not theoretically important to understanding what it is to be in certain mental states (Clark, chapter 1).� The major objection to this theory is that without signi�cant intentional states, we cannot form a computational model. We need to treat the brain as more than a black box which takes stimuli as input and produces actions as output. On the other end of the scale are identity theories, which attempt to link directly mental states and physiological states. This theory fails to explain the possibility of intelligent beings who have a biology entirely different than ours. In the middle is Turing-machine functionalism. Clark notes that �What is essential to being in a certain mental state... is being in a certain abstract functional state.� These abstract functional states are linked to physiological states, but this linkage is only an artifact of our bodies, it is not important to the understanding of mental states.

Clark describes an objection to functionalism raised by Block. Block claims that by abstracting out the physical instantiation of the mind, you can arrive at systems, such as the Chinese message passing system, which could appear to have mental states, but which intuitively lack whatever it is which makes something �intelligent�. I think that this argument is �awed. Given suf�cient structure and complexity, a system should be capable of having mental states, regardless of how oddly constructed it may seem at �rst glance. The cognitive signi�cance of the state of a Turing machine, whether instantiated in human brain tissue, silicon chips, or rubber bands and chewing gum, remains the same.

This is what makes Turing-machine functionalism so attractive.

Another disagreement I have with Block is his notion of what makes a mental state intentional. He claims that �Intentional states represent the world as being a certain way.� He gives the example of tree rings having intentional states, representing the age of the tree. But, he then gives the example of the �swamp replica� of an individual, which (he claims) possesses no intentional states. I fail to see why the representations of the world encoded in the swamp being are any different from the representations of the world in the original person. Block thinks they are quite different.

[The swamp creature’s] utterance is not in any way causally affected by Gorbachev or Thatcher or England, or by Gorbachevian or Thatcherian or English states of the world, so how can it be regarded as being about Gorbachev or Thatcher or England? The swamp creature is simply mouthing words. Hamlet, if written by an in�nite number of monkeys pounding on typewriters, rather

than a single British playwright expounding on human nature, would be no less a great work of literature.

Block also says that �Pain, for example, would seem to be a mental state that has no intentional content.� I completely disagree. Being in pain is a mental state closely associated with certain physical states in the world. For humans, this means having certain nerves and receptors stimulated. Even creatures which don’t have the same kind of pain receptors we do would probably react in similar ways to pain, leading me to believe that being in pain is an abstract intentional mental state, just the same as believing the earth is round.

In general, I do believe that Turing-machine functionalism is the proper theory of mental states. It abstracts cognitive processes away from their medium, so it represents all possible kinds of cognition, and at the same time, allows analysis of how stimuli, beliefs, desires, and actions are related.